OpenStudy (anonymous):

Which equation represents the axis of symmetry of the function y = –2x² + 4x –6? y = 1 x = 1 x = 3 x = –3

OpenStudy (accessdenied):

the axis of symmetry would be the x-value for the vertex (since that's also where the graph stops going one way and starts going the other way.) for y= ax^2 + bx + x axis of symmetry: x = -b / 2a

OpenStudy (accessdenied):

+bx + c, not bx + x.

OpenStudy (anonymous):

x = 1; x= -b/2a = -4/(2 * -2) = -4/-4 = 4/4 = 1/1 = 1